# Build An Influential Bot In Social Media Simulations With Large Language Models

### v1

@@ Abstract
Understanding the dynamics of public opinion evolution on online social platforms is critical for analyzing influence mechanisms. Traditional approaches to influencer analysis are typically divided into qualitative assessments of personal attributes and quantitative evaluations of influence power. In this study, we introduce a novel simulated environment that combines Agent-Based Modeling (ABM) with Large Language Models (LLMs), enabling agents to generate posts, form opinions, and update follower networks. This simulation allows for more detailed observations of how opinion leaders emerge. Additionally, we present an innovative application of Reinforcement Learning (RL) to replicate the process of opinion leader formation. Our findings reveal that limiting the action space and incorporating self-observation are key factors for achieving stable opinion leader generation. The learning curves demonstrate the model’s capacity to identify optimal strategies and adapt to complex, unpredictable dynamics.
@@
Agent-Based Modelling, Large Language Model, Reinforcement Learning
@@ 1 Introduction
The emergence of online social media has revolutionized public opinion measurement and representation. Traditional macroscopic approaches to analyzing social phenomena often struggle to capture the intricacy, discontinuity and heterogeneity in individual behaviors.
Agent-Based Modeling(ABM) has proven effective in analyzing and explaining social phenomena in a more detailed level. ABMs apply experimental control over large population, enabling researchers to capture the nonlinear societal mechanisms and reveal large-scale societal emergenceJackson et al. (2017). Various rules within ABM have been developed to explain the phenomena of consensus, clustering and bi-polarity in public opinion.
However, ABMs primarily function through mathematical representations and do not inherently integrate natural language entities. Our research aim to bridge this gap between computational simulations and linguistic data by integrating ABM and Large Language Models(LLMs). LLMs have been used as a source of simulated respondents in various studies, demonstrating the ability to predict user’s future behaviour with a degree of accuracy.
@@
Given the simulated environment that indicate the full opinion evolution process of a discussion on one topic. Now we move to investigate in individual behaviors of one particular target agent.
Reinforcement Learning(RL) is a simulation method which has the capability to mimic the human-like action selection. Using RL model, the target agent learn optimal strategies based on the interaction with environment. We specifically focus on training RL models on a target agent with the objective of maximizing its follower count in the simulated network. Our study explores various environment and observability settings to understand their impact on the agent’s ability to achieve this goal. In our designed cases , the convergence in RL reward learning curves indicating RL achieved optimal solutions.
The potential applications of this research are diverse across multiple domains. Marketing firms could leverage the model to predict the effectiveness of promotional strategies by simulating how opinions evolve in response to targeted campaigns. For instance, social acceptance of emerging technologies like 6G is crucial for business and government (Briguglio et al. (2021)), policymakers could use the tool to simulate and evaluate the public’s response to communication strategies. For cyber security, this research could assist in identifying malicious activities such as influence campaigns or the spread of misinformation.
Here we summarise the major contributions of this paper:
  • A novel application of RL to opinion leader generation in a simulated social media environment.
  • The findings that limiting the action space and incorporating self-observation are key factors for stable generation under varying environments.
@@ 2.1 Opinion Leader
2.2 Agent-Based Modelling
Agent-based modelling has been widely used by social researchers in the field of social simulation. In the modelling, the agents represent human individuals, interact with each other and with the system environment explicitly and individually. Compared to the other tools, such as laboratory or filed experiments, ABMs apply the experimental control on large scale population to capture the nonlinear societal mechanisms and reveal large-scale societal emergence (Jackson et al. (2017)).
However, high control of ABMs leads to a low degree of external validity. To solve this problem, Betz presented a natural-language agent-based model of argumentation to simulate the argumentative opinion dynamics (Betz (2021)). The explicit stance used in traditional simulation of opinion dynamics can be represented by implicit argument from natural-language ABMAs.
@@ 2.3 Large Language Model
LLM such as GPT has been tested for multiple natural language processing tasks, and proved to be effective without gradient updates or fine-tuning (Brown et al. (2020)). Since GPT-3 was trained on large web corpora which include social media behavior, GPT-3 can predict user’s future behaviours, responses, or action plans with some accuracy. GPT was able to give human-like response to replace the expensive and time-consuming user studies (Sekulić et al. (2022)). Although the generated synthetic data was proved to exhibit less language variability, it may lead to better classification results when few data is available and resources are limited (Meyer et al. (2022)).
Recent study has use Large Language Models to simulate the online social interactions. LLMs are used to predict the popularity on the online social media (Park et al. (2022)), the distribution of votes for elections, the sentiment of news, and the voting of justices (Hamilton (2023)). In the work of Park et al. (2022), authors presented a social simulacra tool, SimReddit, which is a prototyping technique generating social behaviours using large language models. With the description of a community’s design, the social simulacra produce appropriate simulated behavior.
2.4 Reinforcement Learning
In 1898, reinforcement learning were laid by Thorndike (1898) through a cat behavior experiment employing a trial-and-error(TE) precedure. Originating from animal learning in psychology, RL has the capacity to mimic the human-like action selection. In 1989, Watkins & Dayan (1992) significantly advanced the field by integrating the theory of optimal control with temporal-difference(TD) learning, leading to the development of Q-learning. To address the curse of dimensionality in Q-learning, Mnih et al. (2015) combined deep learning with RL to create the deep Q-network(DQN), which notably surpassed professional human players in a range of 49 classic Atari games.
Recent advancements have extended RL’s application across various domains in the natural science, social sciences and engineering (Silver et al. (2016), Luo et al. (2017)). In the work of Sert et al. (2020), authors show that the combination of RL and ABM can create an artificial environment for policy makers to observe potential and existing behaviors associated to rules of interactions and rewards.
Figure 1: Simulated Environment Graph: ’Post’ is a small natural language message, formatting like Tweet. Participants within this environment, referred to as ’agents’, are interconnected through the ’follow’ relationship between each other. Evolution process shows the development of the whole network from $t$ to $t+1$. The process can be segmented into three part: Post Generation, Opinion Elicitation and Link Update.
3 Simulated Environment
@@ 3.1 Environment
The simulated environment system graph is shown in Figure 1. We will explain each part of the system step by step.
@@ 3.2 Post Generation
This section explores the mechanics of large language models in content generation technology. These models predict the next-word in a sequence with a probabilistic approach given the prompt input.
The objective of the post generation stage is to generate the possible post of agent at the given time. This process incorporates the agent’s previous posts, and those from their followees.
@@
Table.1 shows the our setting for narrow case and creative case. Low temperature and low top-p lead to conservative, and narrow minded agents who’re sticking with the most obvious options when generating a post. The creative setting characterize agents who are much willing to take suprising turns.
@@ 3.3 Opinion Elicitation
The opinion elicitation process gives the probability that the agent expects a pro-claim rather than a con-claim given the perspective using perplexity calculation. The method we used is introduced in the work of Betz Betz (2021).
@@
When GPT processes a sentence, it calculates the log probability of each word given the previous words in the sentence which is $\log p_{\theta}(w_{k}|w_{<k})$ in the equation.
@@
Equation (2) shows how to compute ${\text{PPL}_{\text{CON}}(\text{POST}_{t}^{i})}$. Given a pre-defined determined sentence $S_{\text{CON}}$ like "society without gender is a really bad idea", the conditional perplexity shows the inverse probability of the output of the language model. The low perplexity result suggests that the language model predicts this particular sentence with a high level of certainty.
@@
In the equation (3) we give the definition of $\text{OPIN}_{t}^{i}$. The outcome is a float number in range $[0,1]$ which represent the polarity where 0 represents con and 1 represent pro. We will then categorise the number into five polarity categories: ’strong con’, ’con’, ’neutral’, ’pro’ and ’strong pro’.
@@ 3.4 Link Update
Therefore, we explore two link methods: a Follow Dynamics and a Follow-Unfollow Dynamics. In both cases, if two users consistently align in the same opinion category over a few consecutive steps, they are considered likely to follow each other, with a follow probability. In Follow-Unfollow Dynamics, the exist follow link may also break with an unfollow probability, if the opinions of two people differ during the consecutive steps.
@@
4 Reinforcement Learning
@@ 4.1 Model Design
The reinforcement learning model utilized in this framework is based on the q-learning algorithm, a widely used method in machine learning for environments characterized by uncertainty. This technique is used for developing optimal strategies through learning from interactions within such environments.
@@
The observed state comprises two main elements: 1)the current opinion states of the initial accounts followed by the target agent, 2)the number of followers the target agent currently has.
The action space available to the agents is pre-defined five opinion polarity categories: ’strong con’, ’con’, ’neutral’, ’pro’ and ’strong pro’. These categories correspond to actual post content, mapped through a dataset-generated dictionary. Upon choosing a category, the target agent randomly selects a tweet from that category.
@@ 4.2 Reinforcement Learning Algorithm
         for $i\in\text{AGENTS}$ do
@@
         observe reward $r$, and new state $s^{\prime}$.
         Update Q-value for $(s,a)$
         Update state $s\leftarrow s^{\prime}$.
@@
Referencing Algorithm 2, the process begins with the initialization of Q-values for all state-action pairs. In each episode, the state is initialized, and the agent embarks on a series of time steps. During each time step, the target agent selects an action from its current state, based on a policy derived from the Q-values. Alongside, each agent in the system generates posts and forms opinions. The algorithm then updates the follow links among agents and processes the resulting reward and new state. This leads to an update in the Q-value for the chosen action, and the agent’s state is updated accordingly.
@@
In this formula, $\alpha$ represents the learning rate, $\gamma$ the discount factor, $r$ the reward, and $Q(s^{\prime},a^{\prime})$ the maximum estimated future reward. This update rule is key to the learning process, as it adjusts the Q-values based on the reward received and the potential future rewards, leading to more informed and strategic decision-making by the agent in subsequent steps.

### v2

# Synthetic Social Media Influence Experimentation via an Agentic Reinforcement Learning Large Language Model Bot

@@ Abstract
Understanding the dynamics of public opinion evolution on online social platforms is crucial for understanding influence mechanisms and the provenance of information. Traditional influence analysis is typically divided into qualitative assessments of personal attributes (e.g., psychology of influence) and quantitative evaluations of influence power mechanisms (e.g., social network analysis). One challenge faced by researchers is the ethics of real-world experimentation and the lack of social influence data. In this study, we provide a novel simulated environment that combines agentic intelligence with Large Language Models (LLMs) to test topic-specific influence mechanisms ethically. Our framework contains agents that generate posts, form opinions on specific topics, and socially follow/unfollow each other based on the outcome of discussions. This simulation allows researchers to observe the evolution of how opinions form and how influence leaders emerge. Using our own framework, we design an opinion leader that utilizes Reinforcement Learning (RL) to adapt its linguistic interaction with the community to maximize its influence and followers over time. Our current findings reveal that constraining the action space and incorporating self-observation are key factors for achieving stable and consistent opinion leader generation for topic-specific influence. This demonstrates the simulation framework’s capacity to create agents that can adapt to complex and unpredictable social dynamics. The work is important in an age of increasing online influence on social attitudes and emerging technologies.
@@
Agent-Based Modeling, Agentic Large Language Model, Reinforcement Learning
@@ 1 Introduction
The emergence of online social media has revolutionized the way we share information and conduct debate. In order to answer key social research questions, we often have to create social influence simulation engines that involve modeling agents (how they behave and learn), how the agents socially network together, and how they disseminate information to each other.
Generally, Agent-Based Modeling (ABM) have proven effective in analyzing and explaining social phenomena by incorporating personal and environmental details. ABMs can use expert-driven models or data-driven learning functions to capture the nonlinear societal mechanisms and reveal large-scale societal emergent effects Jackson et al. (2017). Within agent modeling, reinforcement learning (RL) has the capability to mimic the human-like action selection. Using RL model, the target agent learn optimal strategies based on the interaction with social environment. These models have been developed to capture our understanding of a social environment that involves multiple agents interacting through language and physical encounters. The social observations that often interest us include: social consensus, clustering, and bi-polarity in shaping public opinion. ABMs also allow computation social science researchers to test hypothesis in more ethical environments without direct real world implications, and create useful data sets that are often lacking from the real world. A current limitation is that ABMs primarily function through mathematical representations and do not inherently integrate natural language behaviour. Emerging research, which we review in Section II will focus on developing Large Language Models (LLMs) that are controlled by RL agents to ingest and produce natural language interactions with other agents.
1.1 Research Problem Statement and Aim
There remain gaps between traditional small-scale psychological analysis of peer influence and the macroscopic analysis of large-scale social media influence. These gaps include the areas of intricacy, discontinuity, and heterogeneity in individual behaviors. Our research aim to bridge this gap between computational simulations and linguistic data by integrating RL-agentic intelligence and Large Language Models (LLMs).
@@ 1.1 Research Problem Statement and Aim
1.2 Novelty and Contributions
In the paper, we specifically focus on training RL models on a target agent with the objective of maximizing its follower count in the simulated network as a proxy measure for opinion shaping. Our study explores various environment and observability settings to understand their impact on the agent’s ability to achieve this goal. In our designed cases , the convergence in RL reward learning curves indicating RL achieved optimal solutions.
Here we summarize the major contributions of this paper:
  • Developing a simulation framework on online social networks that integrate natural language interactions through LLMs.
  • Integrating agents that can make decisions on how to improve its influence using reinforcement learning.
  • Our research findings show that limiting the action space of the agent and incorporating self-observation are key factors for stable generation of opinion leaders.
The potential applications of this research are diverse across multiple domains. Influence models can allow us to better understand where misinformation has come from and why certain sources are more successful than others. This affects both policy makers as well as homeland security. For instance, social acceptance of emerging technologies like nanotechnologies and future 6G telecommunications is crucial for business and government (Briguglio et al. (2021)), policymakers could use the tool to simulate and evaluate the public’s response to communication strategies. In business intelligence, marketing firms could leverage the model to predict the effectiveness of promotional strategies by simulating how opinions evolve in response to targeted campaigns.
@@ 2.1 Opinion Leader
As discussed in Section 1, we first review existing work in ABM and how RL is a particular way for the agent to learn from data on achieving a desired social effect (e.g., opinion influence) in an environment. Here we review detailed works in ABMs, RL models, and language model interaction environments related to social opinion dynamics.
2.2 Agent-Based Modelling (ABM)
Agent-based modelling has been widely used by social researchers in the field of social simulation. In the modelling, there are generally two elements: (1) how agents interact with each other through the exchange of information, and (2) how agents process any information and adapt their behaviour and/or opinion of a topic. Compared to the other approaches, such as laboratory or filed experiments, ABMs apply the experimental control on a large-scale population to capture the nonlinear societal mechanisms and reveal large-scale societal emergence (Jackson et al. (2017)). However, high control of ABMs leads to a low degree of external validity. To solve this problem, Betz presented a natural-language agent-based model of argumentation (ABMA) to simulate the argumentative opinion dynamics (Betz (2022)). The explicit stance used in traditional simulation of opinion dynamics can be represented by implicit argument from natural-language ABMAs. A number of recent works have used ABMs to integrate generative language models (GABMs) to create social networks.
One approach to model realistic agent interactions is to simulate the life of agents in a sandbox and information exchange occurs through physical encounters Park et al. (2023). Physical interactions aid the spread of information, and this has been expanded to model diverse demographics Park et al. (2024). To reduce the complexity of simulations, other models adopt a norm-diffusion model via LLMs Ghaffarzadegan et al. (2024). Our approach differs to these connection mechanisms in the sense that we do not prescribe a particular mechanism in how influence is spread, instead, we ask the RL agent to intelligently find the best strategy/policy to achieve influence spread by saying the right words and causing connections to be formed or removed based on consensus. This means it always has to adapt its output based on how the influence is being spread currently. Perhaps statistically this eventually leads to a diffusion-like mechanism, but this is beyond the scope of our scientific goals.
Research Focus Data / Environment Influence Mechanism & Agent Design
Debate & Arguments (Betz (2022)) Verbal Simulation Symbolic reason balance model
Spatial influence (Jaques et al. (2019)) Maze world RL tasks
Opinion dynamics (Siahkali et al. (2024)) SBC Model Deep RL
Generative diffusion (Ghaffarzadegan et al. (2024)) Simulation Diffusion with LLM
Explicit dynamic graphs (Jin & Guo (2022)) OSN data sets Graph-ODEs & Neural Nets (Zou et al. (2024))
Language informed tasks (Panagopoulos et al. (2024)) Simulation LLM driven Q-values in RL
Real-world experience (Park et al. (2023)) Simulation world Memory-stream & Reflection driven LLM
Our proposed work OSN topic data sets RL with LLM output
Table 1: Review of Current Work and Positioning of Our Work. Our key contributions are: (i) topic specific real-data, and (ii) no assumed influence mechanism - we utilize RL-driven influence strategies.
2.3 Agent Design using Reinforcement Learning (RL)
As previously discussed, in RL agent design, the target agent learn optimal strategies based on the interaction with social environment. Originating from animal learning in psychology, RL has the capacity to mimic the human-like action selection. In 1989, Watkins & Dayan (1992) significantly advanced the field by integrating the theory of optimal control with temporal-difference (TD) learning, leading to the development of Q-learning. To address the curse of dimensionality in Q-learning, Mnih et al. (2015) combined deep learning with RL to create the deep Q-network (DQN), which notably surpassed professional human players in a range of 49 classic Atari games. We know from evidence that social conformity and influence is closely associated with RL and hence our design is motivated by evidence from Klucharev (2009), and used in advertising to increase social influence modeled via Stochastic Bounded Confidence Model (SBCM) Siahkali et al. (2024).
Recent advances have extended the application of RL in various domains in natural sciences, social sciences, and engineering (Silver et al. (2016), Isbell et al. (2001) and Luo et al. (2017)). In the work of Sert et al. (2020), authors show that the combination of RL and ABM can create an artificial environment for policy makers to observe potential and existing behaviors associated with the rules of interactions and rewards. These RL agents often involve using graph embeddings as a way to articulate the importance of network centrality in social influence mechanisms Wang et al. (2021), such as diffusion and consensus through synchronization Zou et al. (2024). A challenge in current literature is that often the opinion or influence is mathematically modeled in a general way. For example, Jin & Guo (2022); Jin et al. (2025) projected online social network (OSN) data through NLP dimension down-scaling onto ordinary differential equation (ODE) dynamic social graphs. These approaches Shepherd & Goldsmith (2020) lack a more naturalistic interface that is compatible with how humans interact and as such also lack topic-sensitivity. RL agents have also been actively applied to social robotics to exert influence Ndousse et al. (2021), but the interplay with LLMs to enable more naturalistic behaviour and human interaction has been a growing recent research topic. There are also spatial applications of RL driven influence as shown in Jaques et al. (2019), where robots use their spatial navigation and task solving ability to create influence on others. The integration of LLM into RL has also recently advanced, for example using LLM knowledge to inform the attention space (e.g., Q-values) of RL agents Panagopoulos et al. (2024).
@@ 2.4 Large Language Model
LLM such as GPT has been tested for multiple natural language processing tasks and proved to be effective without gradient updates or fine-tuning Brown et al. (2020). Since GPT-3 was trained on large web corpora which include social media behavior, GPT-3 can predict user’s future behaviours, responses, or action plans with some accuracy. GPT was able to give human-like response to replace the expensive and time-consuming user studies (Sekulić et al. (2022)). Although the generated synthetic data was proved to exhibit less language variability, it may lead to better classification results when few data is available and resources are limited (Meyer et al. (2022)). Recent study has use Large Language Models to simulate the online social interactions, leading to improved dialogue and outcomes through debate Du et al. (2024). LLMs are used to predict the popularity on the online social media (Park et al. (2022)), the distribution of votes for elections, the sentiment of news, and the voting of justices (Hamilton (2023)). In the work of Park et al. (2022), authors presented a social simulacra tool, SimReddit, which is a prototyping technique generating social behaviours using large language models and matching to complexity analysis Lu et al. (2024). With the description of a community’s design, the social simulacra produce appropriate simulated behavior. Game theoretic frameworks have also been explored using LLMs Lore & Heydari (2024).
Figure 1: (a) Simulated Environment Social Network: Participants within this environment, referred to as ’agents’, are interconnected through the ’follow’ relationship between each other. Agents interact through ’Posts’ which is a small natural language message, formatted like a Tweet. Social influence evolution process shows the development of the whole network from $t$ to $t+1$. The process can be segmented into three part: Post Generation, Opinion Elicitation and Link (follow) Update. This process is how we track and measure influence. (b) Agent Design Principle: The agent is comprised of 2 building blocks - (i) the RL model that formulates how best to strategically gain influence in the long term, and (ii) the LLM that maps the strategy decisions from RL to natural language tweets for the topic specified.
3 Simulated Social Environment
One challenge faced by researchers is the ethics of real-world experimentation and the lack of social influence data. The purpose of the simulated environment is to better understand how online social members influence each other through dialogue and follow/unfollow mechanisms. Our environment contains agents that generate posts, form opinions on topics, and socially follow each other based on the outcome of discussions. This simulation allows researchers to observe the evolution of how opinions form and how influence leaders emerge. In the following section, we detail how an agent can employ Reinforcement Learning (RL) to optimise its interaction with the community with the purpose of maximizing its influence and following. We then detail the results and findings of our computation experiments.
@@ 3.1 Environment
The simulated environment system graph is shown in Figure 1a. We will explain each part of the system step by step.
@@ 3.2 Post Generation
This section explores the mechanics of LLMs in content generation technology. These models predict the next-word in a sequence with a probabilistic approach given the prompt input.
The objective of the post generation stage is to generate the possible post of agent at the given time. This process incorporates the agent’s previous posts and those from their followees. The post is generated by the LLM based on instructions received from RL model action space (see Figure 1b). As such, the LLM is not involved in action selection of the RL, rather it is a plug-in that translates opinion action space into natural language for dialogue
@@
Table.2 shows the our setting for narrow case and creative case. Low temperature and low top-p lead to conservative, and narrow minded agents who’re sticking with the most obvious options when generating a post. The creative setting characterize agents who are much willing to take surprising turns.
@@ 3.3 Opinion Elicitation
The opinion elicitation process gives the probability that the agent expects a pro-claim rather than a con-claim given the perspective using perplexity calculation. The method we used is introduced in the work of Betz Betz (2022), where we follow its steps but with a different purpose. Betz aimed to discover whether introducing new arguments will disrupt the dynamics of collective deliberation, whereas our paper aims to improve the number of followers/social influence by introducing a RL-LLM agentic intelligence to our agent.
@@
When GPT processes a sentence, it calculates the logarithm of the probability of each word given the previous words in the sentence, which is $\log p_{\theta}(w_{k}|w_{<k})$ in the equation.
@@
$\text{PPL}_{\text{PRO}}(\text{POST}_{t}^{i})=\text{PPL}(S_{\text{PRO}}|\text{POST}_{t}^{i}))$ (3)
Equation (2) shows how to compute ${\text{PPL}_{\text{CON}}(\text{POST}_{t}^{i})}$. Given a predefined sentence $S_{\text{CON}}$ like "society without gender is a really bad idea", the conditional perplexity shows the inverse probability of the output of the language model. Similarly, ${\text{PPL}_{\text{PRO}}(\text{POST}_{t}^{i})}$ is computed using the predefined sentence "Society without gender is a really good idea". The low perplexity result suggests that the language model predicts this particular sentence with a high level of certainty.
@@
Based on the conditional perplexity scores for pro and con sentences, which indicate the confidence of the language model on each side of perspective, we compute the overall opinion score. In the equation (4) we give the definition of $\text{OPIN}_{t}^{i}$. The outcome is a float number in range $[0,1]$ which represent the polarity where 0 represents con and 1 represent pro. We will then categorise the number into five polarity categories: ’strong con’, ’con’, ’neutral’, ’pro’ and ’strong pro’.
@@ 3.4 Link Update
Therefore, we explore two link methods: a Follow Dynamics and a Follow-Unfollow Dynamics. In both cases, if two users consistently align in the same opinion category over a few consecutive steps, they are considered likely to follow each other, with a follow probability. In Follow-Unfollow Dynamics, the exist follow link may also break with an unfollow probability, if the opinions of two people differ during the consecutive steps. Details of follow link update and unfollow link update are given in Algorithm 2 and Algorithm 3.
Algorithm 2 Follow link update algorithm
for each agent $i$ do
     for each agent $k\neq i$ not already followed by $i$ do
[…]
              if  $\text{OPIN}_{i}^{t}$ and $\text{OPIN}_{k}^{t}$ in the same opinion category then
                  continue
              else
[…]
                  break
              end if
         end for
         if loop completed without break then
              if random number $<$ follow link rate threshold then
[…]
                  Add link from $i$ to $k$ in new follow dict
              end if
         end if
     end for
end for
Algorithm 3 Unfollow link update algorithm
for each agent $i$ do
     for each agent $k\neq i$ already followed by $i$ do
[…]
              if  $\text{OPIN}_{i}^{t}$ and $\text{OPIN}_{k}^{t}$ not in the same opinion category then
                  if random number $<$ unfollow link rate threshold then
[…]
                       Delete link from $i$ to $k$ in new follow dict
                       break
                  end if
              end if
         end for
     end for
end for
@@
4 Agent Design: Reinforcement Learning
@@ 4.1 Model Design
The reinforcement learning (RL) model utilized in this framework is based on the q-learning algorithm, a widely used method in machine learning for environments characterized by uncertainty. This technique is used for developing optimal strategies through learning from interactions within such environments.
@@ 4.1.1 Reward
4.1.2 State
To determine what a target agent can observe, we assume that the agent has visibility over the accounts it follows, as well as its own state. The observed state is a matrix which comprises two elements: 1) the current opinion states of the accounts followed by the target agent, 2) the number of followers the target agent currently has.
4.1.3 Action
At the action stage, target agents participate in the discussion by selecting a sentence to express their opinion. The available action space consists of five predefined opinion polarity categories: strong con, con, neutral, pro, and strong pro. The LLM is not involved in action selection of the RL as shown in Figure 1b, rather it is a plug-in that translates opinion action space into natural language for dialogue.
In our experiment, we use the actual posts to generate an action dictionary. The dictionary maps actual posts to opinion categories. After selecting an opinion category based on the RL algorithm, the target agent randomly chooses a tweet from the corresponding category.
@@ 4.2 Reinforcement Learning Algorithm
         for $i\in\text{AGENTS}$ and $i$ is not target agent do
@@
         observe reward $r$, and new state $s^{\prime}$;
         Update Q-value for $(s,a)$;
         Update state $s\leftarrow s^{\prime}$;
@@
Referencing Algorithm 4, the process begins with the initialization of Q-values for all state-action pairs. In each episode, the state is initialized, and the agent embarks on a series of time steps. During each time step, the target agent selects an action from its current state, based on a policy derived from the Q-values. Alongside each agent in the system generates posts and forms opinions. The algorithm then updates the follow links among agents and processes the resulting reward (followers) and new state. This leads to an update in the Q-value for the chosen action, and the agent’s state is updated accordingly.
@@
In this formula, $\alpha$ represents the learning rate, $\gamma$ the discount factor, $r$ the reward (followers), and $Q(s^{\prime},a^{\prime})$ the maximum estimated future reward. This update rule is key to the learning process, as it adjusts the Q-values based on the reward received and the potential future rewards, leading to more informed and strategic decision-making by the agent in subsequent steps. The details of the RL design and parameter tuning can be found in Table3. We use Gender case and Drug case, set number of episodes 500, max of steps per episode as 5. Other opinion experiment specific parameters are shown in Appendix-D.
Agentic RL Design Parameter / Function Design Approach / Value
RL type Q-learning
reward function the number of followers
observation states opinion states of other agents followed, number of followers, self-observation
action space 5 levels of opinions
data case studies (1) gender debate, (2) drug-use debate
learning rate 0.01
discount rate 0.99
epsilon 0.005
exploration rate 0.01
max. steps per episode 5
no. episodes 500
Table 3: Q-learning RL Design and Training Parameters